OpenStudy (caerus):

can anyone can integrate this for me:

OpenStudy (caerus):

\[\int\limits_{}^{}\left( \frac{ 2-v }{ v } \right)dv\]

OpenStudy (caerus):

@UnkleRhaukus

OpenStudy (caerus):

2lnv-v+c is right?

OpenStudy (jiteshmeghwal9):

yes

OpenStudy (danjs):

you can rewrite that as two fractions instead of the 1. \[\int\limits[ \frac{ 2 }{ v }-\frac{ v }{ v }]dv = 2 \int\limits \frac{ 1 }{ v }dv - \int\limits dv\]

OpenStudy (caerus):

if e^(2lng-v) hows the IF?

OpenStudy (caerus):

i will get v^2 /e^v ?

OpenStudy (jiteshmeghwal9):

\[\Large e^{2lnv-v}=\frac{e^{lnv^2}}{e^v}=\frac{v^2}{e^v}\]

OpenStudy (caerus):

@jiteshmeghwal9 can you help me with another 10 problems ? ^.^